OpenStudy (anonymous):

The given equation involves a power of the variable. Find all real solutions of the equation. 8x^2-64=0

OpenStudy (slaaibak):

8(x^2 - 8)=0 x=2sqrt(2) or x = -2sqrt(2)

OpenStudy (anonymous):

okay. can you show me what you did there? i'm having real trouble understanding this

OpenStudy (slaaibak):

you first take out the common factor, which is 8: 8(x^2 - 8) = 0 then you use the difference of squares principle: 8(x-sqrt(8))(x+sqrt(8))=0 then x = sqrt(8) or x= -sqrt(8) simplified it's x=2sqrt(2) or x = -2sqrt(2)

OpenStudy (anonymous):

\[8x^2-64=0\] add 64 to both sides \[8x^2=64\] divide both sides by 8 \[x^2=8\] take the square root \[x=\pm\sqrt{8}=\pm2\sqrt{2}\]

OpenStudy (anonymous):

i think it's coming to me... i get where you get the common factor, then the difference of squares is giving me trouble... but i see what you are doing...

OpenStudy (anonymous):

and since \[8=4\times 2\] \[\sqrt{8}=\sqrt{4}\times \sqrt{2}=2\sqrt{2}\]

OpenStudy (anonymous):

how do you know the 2(sqrt)2?? that is really throwing me for a loop.

OpenStudy (anonymous):

why isn't the square root of 8 just 2?